I’m looking for some help with a problem I am having with inputs A2, A3, B2, B3, C2, C3, D2, D3. I turned the synth on last night and the LEDs of these inputs appear to have lit up red with nothing patched. Looking at the scope their levels are at the lowest CV input value. When I patch an LFO in to these inputs there is no change in the value. When I patch a dummy cable in the values remain the same too. Inputs A1, B1, C1 and D1 are all working normally.
I’ve recently updated to the latest stable firmware but didn’t notice any problem at the time. To test I downgraded to my previously installed version but no change.
It is connected via i2c to a few other modules. Everything else in my case looks to be ok.
Any help would be greatly appreciated